A GROCER MIXES 6 POUNDS OF COFFEE THAT IS 60% COLUMBIAN WITH ONE THAT 80% COLUMBIAN HOW MANY POUNDS OF THE SECOND TYPE SHOULD BE USED TO OBTAIN A MIXTURE THAT IS 78% COLUMBIAN?
Answer by stanbon(75887)

A GROCER MIXES 6 POUNDS OF COFFEE THAT IS 60% COLUMBIAN WITH ONE THAT 80% COLUMBIAN HOW MANY POUNDS OF THE SECOND TYPE SHOULD BE USED TO OBTAIN A MIXTURE THAT IS 78% COLUMBIAN?
--------------------
Columbian Equation: 0.60*6 + 0.80x = 0.78(6+x)
Multiply thru by 100 to get:
60*6 + 80(x) = 78*6 + 78x
2x = 18*6
x = 9*6 = 54 lbs. (amt. of 80% coffee in the mixture)
===========================
Cheers,
Stan H.